Quarters and nickels, a total of sixteen.
How many of each, without being seen?
The value of two-twenty is right in this cup.
You will find the answer, a certain cheer up.
Make a table, a graph, an equation will do.
Give it a try and learn something new.
It’s a challenge I know, but don’t give up hope.
The answer is obtainable and within your scope.

Answer by fractalier(6550)

Let the number of quarters and nickels be q and n. then we have
q + n = 16 and
25q + 5n = 220 (in cents) ... now subtract 5 times the first equation and get
-(5q + 5n = 80)
----------------
20q = 140
q = 7 quarters
n = 9 nickels